On January 1, Year 1, the City Taxi Company purchased a new taxi cab for $63,000. The cab has an expected salvage value of $20,000. The company estimates that the cab will be driven 200,000 miles over its life. It uses the units-of-production method to determine depreciation expense. The cab was driven 45,000 miles the first year and 75,000 the second year. What would be the depreciation expense reported on the Year 2 income statement and the book value of the taxi, respectively, at the end of Year 2?

$23,625 and $25,200.

$23,625 and $5,200.

$16,125 and $37,200.

$16,125 and $17,200.
Answer #1

Depreciation expense for year 1

= (Cost - Salvage value)/total life * miles driven

= (63000-20000)/200000*45000 = 9675

Depreciation expense for year 2

= (63000-20000)/200000*75000 = 16125

Book value at the end of the year 2 = 63000 - (9675+16125)

= 37200

Option C is the answer